(3 points) Suppose researchers studying the effect of antibiotic treatment for acute sinusitis compared to symptomatic treatments randomly assigned 156 adults diagnosed with acute sinusitis to one of two groups: treatment or control. Study participants received either a 10-day course of amoxicillin (an antibiotic) or a placebo similar in appearance and taste. The placebo consisted of symptomatic treatments such as acetaminophen, nasal decongestants, etc. At the end of the 10-day period patients were asked if they experienced significant improvement in symptoms. The distribution of responses are summarized below Round all answers to four decimal places. significant Improvement Treatment 54 22 76 Control 58 22 80 Total 112 44 156 What percent of patients in the treatment group experienced a significant improvement in symptoms? What percent of patients in the control group experienced a significant improvement in symptoms? Based on the results of this table, which treatment appears to be more effective for sinusitis A. Control B. Treatment о с. They are approximately equally effective
Answer #1

Percentage of patients in the treatment group experienced a significant improvement = 54/76 = 71%

Percentage of patients in the control group experienced a significant improvement = 58/80 = 72.5%

Based on the results, Control treatment appears to be more effective.